Suppose you just received a shipment of eight televisions. Two of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
Answer by stanbon(75887)

Suppose you just received a shipment of eight televisions. Two of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
P(both work) = 6C2/10C2 = (6*5)/(10*9) = 30/(10*9) = 1/3
--------
P(a random tv is defective:: 2/8 = 1/4
P(a random tv is not defective:: 3/4
---------------------------------------
P(>=1 of 2 is defective) = 1 - (3/4)^2 = 1-9/16 = 7/16
------------
Cheers,
Stan H.
-----------